# On the Client Preference of LLM Fine-tuning in Federated Learning

### v1

@@ Abstract
Reinforcement learning with human feedback (RLHF) fine-tunes a pretrained large language model (LLM) using preference datasets, enabling the LLM to generate outputs that align with human preferences. Given the sensitive nature of these preference datasets held by various clients, there is a need to implement RLHF within a federated learning (FL) framework, where clients are reluctant to share their data due to privacy concerns. To address this, we introduce a feasible framework in which clients collaboratively train a binary selector with their preference datasets using our proposed FedBis. With a well-trained selector, we can further enhance the LLM that generates human-preferred completions. Meanwhile, we propose a novel algorithm, FedBiscuit, that trains multiple selectors by organizing clients into balanced and disjoint clusters based on their preferences. Compared to the FedBis, FedBiscuit demonstrates superior performance in simulating human preferences for pairwise completions. Our extensive experiments on federated human preference datasets – marking the first benchmark to address heterogeneous data partitioning among clients – demonstrate that FedBiscuit outperforms FedBis and even surpasses traditional centralized training.
On the Client Preference of LLM Fine-tuning in Federated Learning
Feijie Wu1, Xiaoze Liu1, Haoyu Wang1, Xingchen Wang1, Jing Gao1 1Purdue University
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capacities in responding to a wide range of open-ended instructions as professionally as human beings. This achievement is attributed to reinforcement learning with human feedback (RLHF) Ziegler et al. (2019); Christiano et al. (2017); Ouyang et al. (2022), a method that aligns an LLM with human preferences. Specifically, it trains a reward model (a.k.a. preference model) with the help of a preference dataset, which includes the comparisons among various completions of given instructions. Then, it fine-tunes the LLM towards generating completions that closely match human preference, evaluated by the reward model.
While empirical studies have validated the effectiveness of RLHF in enhancing LLM performance, RLHF faces a challenge regarding preference data collection. There are two approaches for constructing preference datasets, namely, human efforts Bai et al. (2022); Ganguli et al. (2022); Stiennon et al. (2020) and ChatGPT generation Dubois et al. (2024). The former gathers the preference data from a team of labelers, who rank the completions of each instruction from best to worst. In contrast, the latter entails a set of instructions together with pairwise completions, with ChatGPT Achiam et al. (2023) tasked with selecting the superior completion for each instruction. As LLMs are deployed to serve diverse clients, a preference gap may occur between clients and labelers/ChatGPT, impeding the LLM’s ability to generate responses that satisfy real clients’ tastes. Therefore, there is a demand for a preference dataset that accurately mirrors clients’ preferences in order to facilitate LLM performance.
One approach to meeting the demand is to collect client preference data and build a huge preference dataset, with which an LLM can be fine-tuned on a central entity (a.k.a. server). It is noticed that this approach has been applied to a recent open project named OASST Köpf et al. (2024). However, this approach may be infeasible because most clients refuse the disclosure of their preference data out of privacy concerns.
Figure 1: An outline of RLHF in federated learning.
To this end, we adopt FedAvg, a conventional federated learning (FL) algorithm Konečnỳ et al. (2016); McMahan et al. (2017) that avoids the collection of clients’ raw data. Specifically, each client trains a reward model with their preference data, and the server aggregates the reward models into a global model. While the process seems effective, we observe the following two limitations during training:
@@
  • Excessive Computation Overhead: Conventionally, the reward model is designed to output a scalar value for a given prompt and completion Stiennon et al. (2020). Its optimization is based on reward differences between preferred and dispreferred completions, i.e., the preferred completions should earn greater rewards than the dispreferred ones. As a result, when the optimization involves a data sample in training the model, it should simultaneously retain two computation graphs in the forward and backward pass, leading to significant computation overhead and intensive GPU requirements.
  • Degraded Performance: Preference data are heterogeneous among the clients in view that the preferences vary among the clients. Consequently, each client trains a reward model towards a local minima, deviating from the global optima and resulting in a longer training time to converge when compared to the centralized training. Moreover, the method is likely to suffer from reward hacking, a phenomenon where the reward model heads to a poor performance after several training rounds Askell et al. (2021); Michaud et al. (2020); Tien et al. (2022); Skalse et al. (2022).
@@
In this paper, we propose to address these two limitations and propose effective and computationally efficient methods for preference collection and subsequent fine-tuning. We start with a solution that addresses the first limitation. The key idea is to train a binary selector, which chooses a superior response between two completions under a given instruction. Compared with preference ranking, the binary selector requires much less computation during training. Casting binary selector training into a federated learning setting, we thus propose the federated binary selector training (FedBis), as depicted in the first stage of Figure 1. Each client independently trains the selector with local preference dataset, and the server aggregates the selectors into a global one and broadcasts it to the clients. Afterwards, we utilize the binary selector to enhance the performance of LLM. Specifically, we assume the server holds a set of instructions, together with pairwise responses generated by an LLM. Then, we build a preference dataset with the help of the binary selector and boost the LLM by means of direct preference optimization (DPO) Rafailov et al. (2023).
@@
To further address the performance deterioration due to preference heterogeneity and reward hacking, we propose a method named FedBis with cluster-wise aggregation (FedBiscuit). This method ensembles multiple binary selectors, each trained by the clients possessing similar preferences. In light of privacy concerns, which prevent explicit sharing of clients’ data, the server intermittently collects client losses on all binary selectors. Subsequently, clients are organized into disjoint clusters, and when comparing two completions, the one selected by the majority of binary selectors is deemed better. The proposed method has two main advantages. Firstly, clients with similar preferences jointly train a binary selector, moderating data heterogeneity and mitigating performance deterioration. Secondly, the method alleviates reward hacking by having numerous binary selectors jointly decide on optimal completions.
@@
  • To the best of our knowledge, this is the first feasible framework to achieve RLHF in FL. In detail, the framework trains binary selector(s) with clients’ local datasets, distills the selector(s) toward an LLM, and boosts LLM performance in the meantime. Under this framework, we introduce two methods, i.e., FedBis and FedBiscuit.
  • Previous works offer a number of human preference datasets, but none of them address the FL setting. This is the first work to discuss the possible data partition approaches to build a heterogeneous human preference dataset. To this end, we introduce a benchmark that includes several human preference datasets suitable for FL.
  • We conduct extensive experiments to demonstrate the performance of the proposed FedBis and FedBiscuit. As expected, FedBiscuit demonstrates superior performance over FedBis and even surpasses traditional centralized training. Meanwhile, we present some insights from the empirical studies.
@@ 2 Related Work
LLM Fine-tuning in FL.
Recent studies have increasingly focused on fine-tuning large language models (LLMs) using federated datasets Sun et al. (2024); Ye et al. (2024); Zhang et al. (2023a); Yi et al. (2023); Zhang et al. (2023b). However, these approaches often suffer from high computation and communication costs due to the necessity of training and synchronizing the model with clients. To mitigate these issues, lightweight methods such as black-box fine-tuning Sun et al. (2023); Lin et al. (2023) and offsite-tuning Wu et al. (2024); Kuang et al. (2023) have emerged. Despite their advancements, these methods primarily focus on fine-tuning LLMs for specific downstream tasks, neglecting user preferences in the generated responses. To address this gap, our work aims to align LLMs with human preferences and introduces a feasible training framework in federated learning.
@@
RLHF typically involves supervised fine-tuning, reward modeling, and reward optimization, initially popularized by Christiano et al. (2017). Proximal Policy Optimization (PPO) Schulman et al. (2017) is a common RLHF algorithm, yet it struggles with instability, inefficiency, and high resource demands Choshen et al. (2019); Engstrom et al. (2020). These challenges have led to the development of alternative methods, such as Direct Preference Optimization (DPO) Rafailov et al. (2023) and others Dong et al. (2023); Zhao et al. (2023); Azar et al. (2024); Ethayarajh et al. (2024); Gulcehre et al. (2023), which offer more stable and efficient solutions. However, these methods typically operate within a centralized training framework, where the LLM owner retains control over the preference data. In contrast, our work seeks to expand data sources and incorporate real user preferences during the fine-tuning of the LLM.
3 FedBis: A Feasible Framework for Achieving RLHF in FL
The objective of RLHF is to align a pretrained language model with human preferences. RLHF comprises two phases: (i) preference modeling and (ii) reinforcement-learning fine-tuning. The first phase aims to develop a model that simulates human preferences to select the superior options from numerous pairwise completions. Subsequently, the second phase enhances the language model’s performance by creating a preference dataset, enabling the model to generate responses preferred by humans. In the following, We describe the proposed FedBis that achieves RLHF in FL in the first two subsections, followed by a brief discussion of its limitations that motivate the proposed FedBiscuit presented in Section 4.
3.1 Preference Modeling
3.1.1 Problem Formulation.
In preference modeling, our objective is to train a binary selector using data from multiple clients. Consider an FL system with $M$ clients, coordinated by a central server. Denote the weight of client $m$ as $p_{m}$ such that $\sum_{m\in[M]}p_{m}=1$, and we aim to optimize the following objectives:
@@ 3.1 Preference Modeling
where $F_{m}(\phi)$ is the expected loss on client $m$ given the binary selector $\phi$. Suppose client $m\in[M]$ holds a set of pairwise data with the size of $n_{m}$, i.e., $\hat{\mathcal{D}}_{m}=\{(x_{i},y_{i,w},y_{i,l})\}_{i\in[n_{m}]}$, where $x_{i}$ is the prompt, $y_{i,w}$ is the preferred completion out of the pair of $y_{i,w}$ and $y_{i,l}$. We reorganize these data and build a preference dataset $\mathcal{D}_{m}$ to be $\{(x_{i},y_{i,w},y_{i,l},0),(x_{i},y_{i,l},y_{i,w},1)|(x_{i},y_{i,w},y_{i,l})\in\hat{\mathcal{D}}_{m}\}$ for training, in which each contains the prompt, a pair of completions and preference selection. Apparently, this dataset eliminates the position effects, and we can train the selector as a classification task. Therefore, we utilize cross-entropy (CE) loss $\ell_{CE}$ to optimize the selector and formulate the expected loss as

### v2

# Towards Federated RLHF with Aggregated Client Preference for LLMs

@@ Abstract
Reinforcement learning with human feedback (RLHF) fine-tunes a pretrained large language model (LLM) using user preference data, enabling it to generate content aligned with human preferences. However, due to privacy concerns, users may be reluctant to share sensitive preference data. To address this, we propose utilizing Federated Learning (FL) techniques, allowing large-scale preference collection from diverse real-world users without requiring them to transmit data to a central server. Our federated RLHF methods (i.e., FedBis and FedBiscuit) encode each client’s preferences into binary selectors and aggregate them to capture common preferences. In particular, FedBiscuit overcomes key challenges, such as preference heterogeneity and reward hacking, through innovative solutions like grouping clients with similar preferences to reduce heterogeneity and using multiple binary selectors to enhance LLM output quality. To evaluate the performance of the proposed methods, we establish the first federated RLHF benchmark with a heterogeneous human preference dataset. Experimental results show that by integrating the LLM with aggregated client preferences, FedBis and FedBiscuit significantly enhance the professionalism and readability of the generated content.
@@ 1 Introduction
Large language models (LLMs) exhibit broad knowledge coverage as they are pretrained on a large corpus. To elevate the quality of their generated content to match the professionalism and readability of human writing, a common approach is to fine-tune these models using reinforcement learning with human feedback (RLHF) Ziegler et al. (2019); Christiano et al. (2017); Ouyang et al. (2022). This process relies on preference datasets, which are constructed through two main methods: human evaluation Bai et al. (2022); Ganguli et al. (2022); Stiennon et al. (2020) and ChatGPT-based ranking Dubois et al. (2024b). In the human-effort approach, a set of instructions (a.k.a. prompts or user queries) is paired with multiple model completions (a.k.a. generated responses), and a team of labelers ranks completions of each instruction from best to worst. In the ChatGPT-based method, the dataset is developed by feeding ChatGPT with multiple inputs encompassing an instruction and a pair of completions, where it selects the superior completion for each input. As LLMs are deployed to serve diverse users, there could be a gap between the preferences of real-world users and those of lablers/ChatGPT, hindering the LLM’s ability to generate responses that align with the users’ preferences.
Figure 1: Comparison between standard and federated RLHF
Therefore, there is a need for a preference dataset that accurately reflects real world users’ preferences in order to enhance the ability of LLM in content generation.
One straightforward approach to meeting this need is to directly collect preference data from a large number of real-world users and build a comprehensive preference dataset, which can be used to fine-tune an LLM on a central server. This strategy has been implemented in recent projects like OASST Köpf et al. (2024). However, this approach is often impractical because it involves collecting LLM users’ inputs and preferences – both highly private and sensitive (shown in the left part of Figure 1). Most users are unwilling to share their data, and such practices may also violate regulations like the GDPR (2016) and CCPA (2023), which prohibit the use of users’ private data for model training without explicit consent.
To address privacy concerns, we propose utilizing Federated Learning (FL) techniques to enable large-scale preference collection from diverse real-world users without requiring them to transmit their preference data to a central server. In our design, we adopt FedAvg, a well-established FL algorithm Konečnỳ et al. (2016); McMahan et al. (2017), to learn user preferences without directly collecting their data. In line with the reward model training used in RLHF Stiennon et al. (2020); Ouyang et al. (2022); Rafailov et al. (2023), each LLM user works as a FL client and locally trains a reward model that can evaluate the quality of model completions. The server then aggregates these client-trained reward models into a global model. While this approach appears effective, we observe the following two limitations during training:
@@
  • Excessive Computation Overhead: Typically, the reward model is a regression model designed to output a scalar value that represents the quality of a model completion Stiennon et al. (2020). Its optimization is based on comparing reward differences between preferred and dispreferred completions—preferred completions should receive higher rewards. However, during optimization, each data sample requires the retention of two computation graphs (one for the preferred and one for the dispreferred completion) throughout the forward and backward passes. This results in significant computational overhead and heavy GPU demands.
  • Preference Heterogeneity and Reward Hacking: Client preferences are heterogeneous, as both instructions and preferences vary across clients. As a result, each client tends to train their reward model toward a local minimum, which deviates from the global optimum, leading to longer convergence times compared to centralized training. Moreover, fine-tuning the pretrained model using the reward model can lead to overfitting, where the proxy reward (the reward model’s output) improves while the actual performance worsens. This phenomenon, known as reward hacking or reward overfitting, has been discovered in several studies Askell et al. (2021); Michaud et al. (2020); Tien et al. (2022); Skalse et al. (2022).
@@
In this paper, we propose to address these two limitations and propose effective and computationally efficient methods for preference collection and subsequent fine-tuning. We start with a solution that addresses computation costs. The key idea is to train a binary selector that identifies the superior response between two model completions. Compared with traditional reward model training, the binary selector requires significantly less computation. Casting binary selector training into a federated learning setting, we develop a federated binary selector training (FedBis) framework, as depicted in the right part of Figure 1. The aggregated binary selector captures the common preferences of a large group of users, and thus can simulate a comprehensive preference dataset, facilitating RLHF approaches (e.g., DPO Rafailov et al. (2023)) to fine-tune the LLM without the need for the real client preference data.
@@
To further address the performance deterioration due to preference heterogeneity and reward hacking, we propose a method named FedBis with cluster-wise aggregation (FedBiscuit). This approach ensembles multiple binary selectors, each trained by clients with similar preferences. Since privacy concerns prevent the explicit sharing of client data, the server intermittently collects the training loss of all binary selectors on the clients. Using this data, clients are grouped into disjoint clusters, and when comparing two completions, the one favored by the majority of selectors is deemed superior. This method has two main advantages: (1) Clients with similar preferences collaboratively train a binary selector, reducing data heterogeneity and improving performance stability, and (2) Reward hacking is mitigated by employing multiple binary selectors, as it becomes difficult for the LLM to generate content that satisfies all binary selectors without making genuine improvements.
@@
  • To the best of our knowledge, this is the first work to employ federated learning technique to enable large-scale user preference collection for RLHF without jeopardizing user privacy. Our proposed federated RLHF model (i.e., FedBis) encodes each client’s preference information into a binary selector and aggregates all clients’ binary selectors to capture their common preferences. By aligning the LLM with the aggregated client preference, we can improve the professionalism and readability of LLM’s generated content.
  • We identify the inherent challenges of federated RLHF, such as preference heterogeneity and reward hacking, and extend FedBis into FedBiscuit with innovative solutions to address these challenges, including grouping the binary selectors of clients with similar preferences to reduce data heterogeneity and employing multiple binary selectors to force the LLM to improve the quality of its generated content.
  • We conduct extensive experiments to evaluate the performance of the proposed FedBis and FedBiscuit. Since no prior work has addressed RLHF in a federated learning setting, we establish the first FL benchmark by creating a heterogeneous human preference dataset. As expected, both FedBis and FedBiscuit show significant performance improvements over the base models, Gemma and LLaMA.
@@ 2 Related Work
Federated Fine-Tuning for LLM.
Recent studies have increasingly focused on fine-tuning large language models (LLMs) using federated datasets Sun et al. (2024); Ye et al. (2024); Zhang et al. (2023a); Yi et al. (2023); Zhang et al. (2023b). However, these approaches often suffer from high computation and communication costs due to the necessity of training and synchronizing the model with clients. To mitigate these issues, lightweight methods such as black-box fine-tuning Sun et al. (2023); Lin et al. (2023) and offsite-tuning Wu et al. (2024a); Kuang et al. (2023) have emerged. Despite their advancements, these methods primarily focus on fine-tuning LLMs for specific downstream tasks, neglecting user preferences in the generated responses. A recent benchmark, OpenFedLLM Ye et al. (2024), introduces FedDPO, which allows federated clients to optimize their local LLMs using DPO loss. While this approach can potentially align LLMs with human preferences, it faces three key challenges: excessive computational overhead, preference heterogeneity, and the risk of reward hacking. To address these limitations, our work aims to enable LLMs alignment with a feasible and sustainable training framework in FL.
@@
RLHF typically involves supervised fine-tuning, reward modeling, and reward optimization, initially proposed by Christiano et al. (2017). Proximal Policy Optimization (PPO) Schulman et al. (2017) is a common RLHF algorithm, yet it struggles with instability, inefficiency, and high resource demands Choshen et al. (2019); Engstrom et al. (2020). These challenges have led to the development of alternative methods, such as Direct Preference Optimization (DPO) Rafailov et al. (2023) and others Dong et al. (2023); Zhao et al. (2023); Azar et al. (2024); Ethayarajh et al. (2024); Gulcehre et al. (2023), which offer more stable and efficient solutions. However, these methods typically operate within a centralized training framework, where the LLM owner retains control over the preference data. In contrast, our work aims to expand data sources and integrate real user preferences without directly collecting their personal data.
3 Preliminary
3.1 Federated Learning (FL)
FL is a distributed training paradigm where a server coordinates various clients toward the same goal, i.e., training a generalized model $\phi\in\mathbb{R}^{d}$ Konečnỳ et al. (2016); McMahan et al. (2017). Consider an FL system with $M$ clients. Denote the weight of client $m$ as $p_{m}$ such that $\sum_{m\in[M]}p_{m}=1$, and we aim to optimize the following objectives:
@@ 3.1 Federated Learning (FL)
where $F_{m}(\phi)$ is the expected loss on client $m$ given the model $\phi$. As a classical FL algorithm, FedAvg can solve the optimization problem by multiple communications between the server and the clients, followed by a number of recent works Wang et al. (2022); He et al. (2023); Fang et al. (2024); Wang et al. (2024a; b); Wu et al. (2024b). In each communication round $r\in[R]$ with the global model $\phi_{r}$, the following steps are conducted:
  • Model broadcast: The server uniformly samples $A$ clients without replacement, denoted by $\mathcal{A}$, broadcasts the global model $\phi_{r}$ to the sampled clients.
  • Local training on client $m\in\mathcal{A}$: The client initializes the local model $\phi_{r,0}^{m}$ with the received $\phi_{r}$. In the next $K$ iterations, the client updates its local model via $\phi_{r,k+1}^{m}=\phi_{r,k}^{m}-\eta\nabla F_{m}(\phi_{r,k}^{m}),k\in[K]$, where $\eta$ is learning rate, and $\nabla F_{m}(\phi_{r,k}^{m})$ is the local gradient on $\phi_{r,k}^{m}$ and estimated by a mini-batch. We denote this $K$-iteration local training by $\phi_{r,K}^{m}=\textsc{Optim}(m,\phi_{r},K)$, which optimizes the model $\phi_{r}$ with the data of client $m$.
  • Global aggregation: The server collects the local model $\phi_{r,K}^{m}$ from clients $m\in\mathcal{A}$ and updates the global model via weighted average aggregation, i.e., $\phi_{r+1}=\frac{M}{A}\sum_{m\in\mathcal{A}}p_{m}\phi_{r,K}^{m}$.
3.2 Reinforcement Learning with Human Feedback (RLHF)
The objective of RLHF is to align a pretrained language model with human preferences so that the model can generate text that is as professional and readable as human writing. DPO Rafailov et al. (2023) is one of the most effective ways to achieve the goal. There is a human-annotated dataset $\mathcal{D}$ comprising multiple samples $(x,y_{0},y_{1},i)$, where $y_{0}$ and $y_{1}$ are two completions under a given instruction $x$, and $i\in\{0,1\}$ indicates that $y_{i}$ is the preferred completion out of the pair of $y_{0}$ and $y_{1}$. Motivated by the Bradley-Terry model Bradley & Terry (1952) on the formulation of human preference distrition $p^{*}(y_{i}\succ y_{1-i}|x)$, DPO aims to optimize the model $\theta$ starting from $\theta_{0}$ via
[…]
4 FedBis: A Vanilla and Feasible Framework for Achieving Federated RLHF
Figure 2: An outline of the proposed FedBis, an RLHF method in federated learning.
We aim to fine-tune an LLM using clients’ preference data, enabling it to generate reader-friendly responses. Since preference data contain sensitive personal information, some clients may be hesitant to share this information due to privacy concerns. Recently, companies have developed on-device pretrained language models (e.g., Phi-3 Abdin et al. (2024) and Qwen Bai et al. (2023)), with the latest iPhone release integrating this technology Gunter et al. (2024). This on-device feature allows clients to ask private and sensitive questions directly on their smartphones, ensuring that even the server (i.e., LLM owner) cannot access the input prompts Wu et al. (2024a).
The proposed FedBis provides a simple yet effective solution through a three-step process to enable model fine-tuning with clients’ preference data while respecting clients’ privacy, as depicted in Figure 2: In the first step, each client builds their own preference dataset, which is originated from the daily queries (a.k.a. prompts or instructions) to a pretrained language model (LM), the model generated a pair of responses to each query, and the client chooses the preferred one. After the construction of the preference dataset, each client independently trains the selector, and the server aggregates the selectors into a global one and broadcasts it to the clients. This communication process repeats for a total of $R$ times, and the clients keep the preference dataset unchanged during the training. Afterward, we utilize the well-trained binary selector to enhance the performance of LLM. Specifically, we assume the server holds a set of instructions, together with pairwise responses generated by an LLM. Then, we build a preference dataset with the help of the binary selector and boost the LLM by means of DPO Rafailov et al. (2023). The rest highlights the key steps of these two phases, while Appendix A offers a detailed description of the proposed FedBis.
Local Training in FL.
@@ 4 FedBis: A Vanilla and Feasible Framework for Achieving Federated RLHF
Suppose client $m\in[M]$ holds a set of pairwise data with the size of $n_{m}$, i.e., $\hat{\mathcal{D}}_{m}=\{(x_{i},y_{i,w},y_{i,l})\}_{i\in[n_{m}]}$, where $x_{i}$ is the prompt, $y_{i,w}$ is the preferred completion out of the pair of $y_{i,w}$ and $y_{i,l}$. We reorganize these data and build a preference dataset $\mathcal{D}_{m}$ to be $\{(x_{i},y_{i,w},y_{i,l},0),(x_{i},y_{i,l},y_{i,w},1)|(x_{i},y_{i,w},y_{i,l})\in\hat{\mathcal{D}}_{m}\}$ for training, in which each contains the prompt, a pair of completions and preference selection. Apparently, this dataset eliminates the position effects, and we can train the selector as a classification task. Therefore, we utilize cross-entropy (CE) loss $\ell_{CE}$ to optimize the selector and formulate the expected loss as